OpenStudy (anonymous):

6+(-6) I'm not sure whether or not it's cancelled out into 0 because a positive plus a negative should equal a negative right? I don't know...

OpenStudy (anonymous):

multiplication + x + =+ + x - = - - x - = + 6+(-6) = 6 - 6

OpenStudy (anonymous):

How is multiplication involved?

OpenStudy (texaschic101):

when there is parenthesis, there is multiplication

OpenStudy (anonymous):

Oooh... I thought it was placed there because of the negative sign... is there an invisible one being multiplied because I doubt anything's being multiplied by 6...

OpenStudy (texaschic101):

this might help...when there is a sign, either positive or negative, before the parenthesis, and no number, you can put in 1. Like this : 6 + 1(-6) 6 - 6 --- because 1 * -6 = -6 answer is 0.

OpenStudy (anonymous):

I see.That makes sense... thank you both.

OpenStudy (texaschic101):

anytime :)